# Compositional Generative Model of Unbounded 4D Cities

### v2

@@ 1 Introduction
To address these problems, we propose CityDreamer4D, a compositional generative model designed for unbounded 4D cities. As shown in Fig. 1, the unbounded 4D city generation framework separates dynamic objects from static scenes. Static scenes are defined by the city layout from Unbounded Layout Generator, arranging elements like roads, highways, vegetation, and buildings, with the capability to extrapolate to unbounded areas. Dynamic objects, such as vehicles, are defined by traffic scenarios generated by Traffic Scenario Generator, which determines their spatial positioning on high-fidelity (HD) maps derived from city layouts. Unlike existing methods that use a single module for all objects, CityDreamer4D divides the generation process into three distinct modules: Building Instance Generator for buildings, Vehicle Instance Generator for vehicles, and City Background Generator for background stuff. These generators leverage a highly compact bird’s-eye-view (BEV) scene representation to ensure efficiency and scalability. The scene parameterization is designed to address the unique characteristics of background stuff and instances: background stuff often features similar appearances with irregular textures, while buildings and vehicles display diverse appearances with regular periodic patterns. To handle these variations, we use generative hash grids for the background and apply periodic positional encodings to each instance. We also place buildings in an object-centric coordinate space and vehicles in an object-canonical coordinate space, using specialized methods designed to capture their compact shapes. Compositor combines the rendered background stuff with the building and vehicle instances to create a unified image.
@@
A preliminary version of this work, named CityDreamer, has been published in CVPR 2024 [34]. We make several extensions in this work compared to the preliminary version. 1) We evolve CityDreamer into CityDreamer4D, enabling 4D city generation through Traffic Scenario Generator and Vehicle Instance Generator, effectively separating dynamic objects from static scenes. 2) We enhance the highly compact BEV representation by incorporating an additional bottom-up height map, enabling the representation of hollow structures in cities, such as highways. 3) We propose Traffic Scenario Generator, which creates high-fidelity maps from city layouts to produce realistic traffic scenarios with vehicles in unbounded cities. 4) We introduce Vehicle Instance Generator, designed to generate vehicle instances within cities using a novel scene parameterization method grounded on the canonical feature space. 5) We build the CityTopia dataset, offering nearly 40k high-quality street-view and drone-view images with both 2D and 3D semantic and instance annotations.
@@ 2 Related Works
Unlike advanced 2D generative models that mainly focus on individual categories or familiar objects, generating scene-level content is more challenging due to the vast diversity and complexity of scenes. Earlier methods [15, 16] generate scenes by synthesizing videos, but they lack 3D awareness and fail to ensure 3D consistency. Semantic image synthesis approaches [24, 62] have shown promising results in generating scene-level content by conditioning on pixel-wise dense correspondences, like semantic segmentation or depth maps. Several techniques [18, 19, 20] generate 3D scenes by performing inpainting and outpainting on RGB images or feature maps, though most can only interpolate or extrapolate a limited distance from the input views and lack true generative capabilities. Significant research has investigated procedural content generation (PCG) for creating natural [63, 21], indoor [64, 65, 66], and urban scenes [67, 68, 23, 69], but the diversity of the generated scenes is limited by the 3D assets used. Recent methods [26, 7, 9] achieve 3D-consistent scenes at an infinite scale through unbounded layout extrapolation. Other approaches [70, 71, 72] focus on indoor scene synthesis, relying on costly 3D datasets [73, 74] or CAD object retrieval [75, 76, 77].
Figure 1: Overview of CityDreamer4D. 4D city generation is divided into static and dynamic scene generation, conditioned on $\mathbf{L}$ and $\mathbf{T}_{t}$, produced by Unbounded Layout Generator and Traffic Scenario Generator, respectively. City Background Generator uses $\mathbf{L}$ to create background images $\mathbf{\hat{I}}_{G}$ for stuff like roads, vegetation, and the sky, while Building Instance Generator renders the buildings $\{\mathbf{\hat{I}}_{B_{i}}\}$ within the city. Using $\mathbf{T}_{t}$, Vehicle Instance Generator generates vehicles $\{\mathbf{\hat{I}}_{V_{i}}^{t}\}$ at time step $t$. Finally, Compositor combines the rendered background, buildings, and vehicles into a unified and coherent image $\mathbf{\hat{I}}_{C}^{t}$. “Gen.”, “Mod.“, “Cond.”, “BG.”, “BLDG.”, and “VEH.” denote “Generation”, “Modulation”, “Condition”, “Background”, “Building”, and “Vehicle”, respectively.
@@ 3.1 Unbounded Layout Generator
where 00 denotes empty spaces in the 3D volumes.
@@ 3.2 Traffic Scenario Generator
where 00 denotes empty spaces in the 3D volumes.
Traffic Scenario Generation. The generation of traffic scenario $\mathcal{T}$ is conceptualized as the frame-by-frame production of semantic maps $\mathbf{S}_{T}=\left\{\mathbf{S}_{T_{t}}\right\}_{t=1}^{n^{T}}$ and height fields $\mathbf{H}_{T}=\left\{\mathbf{H}_{T_{t}}\right\}_{t=1}^{n^{T}}$. To guarantee realistic and continuous placement of dynamic objects, a high-fidelity (HD) map is derived from the city layout $\mathbf{L}$. Unlike the city layout, which only specifies the positions of roads and highways, the HD map includes details about lanes, intersections, and traffic signals. Using the generated HD map, an off-the-shelf model [86] determines the per-frame bounding boxes of dynamic objects. The corresponding semantic map and height field are generated based on the bounding boxes.
@@ 3.3 City Background Generator
[…]
@@ 3.4 Building Instance Generator
Scene Representation. Building Instance Generator also employs the BEV scene representation. It extracts a local window $\mathbf{L}^{B_{i}}$ from the city layout $\mathbf{L}$ with dimensions $N_{B}^{H}\times N_{B}^{W}\times N_{B}^{D}$. This window is centered around the 2D coordinates $(c_{x}^{B_{i}},c_{y}^{B_{i}})$ of the building instance $B_{i}$. The height field and semantic map used to construct $\mathbf{L}^{B_{i}}$ are represented as $\mathbf{H}_{L}^{B_{i}}$ and $\mathbf{S}_{L}^{B_{i}}$, respectively. Since all buildings share the same semantic label in $\mathbf{S}_{L}$, we perform building instantiation by detecting connected components. Notably, real-world building facades and roofs exhibit distinct visual distributions. To capture this, we assign different semantic labels to the facade and roof of each building instance $\rm B_{i}$ in $\mathbf{L}^{B_{i}}$, with the roof assigned to the top-most voxel layer. All other building instances are excluded from $\mathbf{L}^{B_{i}}$ by assigning them a value of 00.
@@
Note that $\mathcal{O}(\cdot)$ is applied separately to each element of the feature $\mathbf{x}$, with the values normalized to the range $[-1,1]$.
@@
[…]
@@
(×$\times 10^{3}$) (km2) Day Night Street Aerial 2D Sem. 2D Inst. 3D Sem. 3D Inst.
@@
CityTopia 3.75 11 36 Synthetic Voxel
@@ 3.5 Vehical Instance Generator
Scene Representation. Vehicle Instance Generator, like Building Instance Generator, leverages the BEV scene representation. It extracts a local window $\mathbf{T}_{t}^{V_{i}}$ from the traffic scenario $\mathbf{T}_{t}$, with dimensions $N_{V}^{H}\times N_{V}^{W}\times N_{V}^{D}$, to generate the vehicle instances within the scene. This window is centered around the 2D coordinates $(c_{x}^{V_{i}},c_{y}^{V_{i}})$ of the vehicle instance $V_{i}$. The height field and semantic map used to construct $\mathbf{T}_{t}^{V_{i}}$ are represented as $\mathbf{H}_{T_{t}}^{V_{i}}$ and $\mathbf{S}_{T_{t}}^{V_{i}}$, respectively. Unlike buildings, vehicle instances are instantiated during the generation of the traffic scenario. Instances other than $V_{i}$ are removed from $\mathbf{T}_{t}^{V_{i}}$ by assigning them a value of 00.
@@
where $\theta\in(-180\degree,180\degree]$ denotes the yaw angle, indicating the vehicle’s heading in the XY-plane relative to the $-y$-axis, while $\gamma\in(-90\degree,90\degree)$ represents the pitch angle, with positive or negative values indicating upward or downward tilt relative to the XY-plane. The feature $\mathbf{f}_{\rm V_{i}}^{(\mathbf{p}^{C},t)}$ corresponding to the vehicle $V_{i}$ at time step $t$ for $\mathbf{p}^{C}$ is derived as
@@ 3.6 Compositor
$\mathbf{I}_{C}=\mathbf{\hat{I}}_{G}\mathbf{M}_{G}+\sum_{i=1}^{n_{B}}{\mathbf{\hat{I}}_{B_{i}}\mathbf{M}_{B_{i}}}+\sum_{i=1}^{n_{V}}{\mathbf{\hat{I}}_{V_{i}}^{t}\mathbf{M}_{V_{i}}^{t}}$ (18)
@@ 4.3 CityTopia Dataset
Virtual City Generation. To build the CityTopia dataset, we design 11 virtual cities in Houdini and Unreal Engine, generating 3D annotations and realistic images with controlled lighting conditions. As illustrated in Fig. 3(a), we use a diverse, high-quality set of approximately 5,000 3D assets from the CitySample project [33] to procedurally generate a city prototype in Houdini [^1] https://www.sidefx.com. This city prototype stores the 6D poses of all 3D assets within the city. Through surface sampling, we can assign each 3D point a semantic and instance label, and by instantiating the city prototype in Unreal Engine [^2] https://www.unrealengine.com, we produce a fully generated virtual city.
Image Collection. Once the virtual city is instantiated in Unreal Engine, camera trajectories are set to generate 3,000 images for cities with buildings and 7,500 for a vehicle-only city. Daytime and nighttime scenes are rendered for each trajectory, with sunlight removed to help the network more easily learn lighting consistency during the generation process. To avoid Moiré effects, each image is sampled 8x spatially and 32x temporally during rendering. As shown in Fig. 3(c), the CityTopia dataset provides a wider range of viewpoints, shown by its broader elevation angles compared to the GoogleEarth dataset, as well as more street-level perspectives, evidenced by the large number of images taken at near-zero altitude.
2D and 3D Annotation. Since the precise 3D annotations are natively generated from the virtual city pipeline, once the camera poses are set in Unreal Engine, 2D annotations are produced by projecting the 3D annotations using the given camera poses. Fig. 3(b) highlights the perfect alignment of 2D and 3D instance annotations with both street-view and aerial-view images in the CityTopia dataset. The last row features a vehicle-only scene, enhancing vehicle generation learning. The accurate vehicle annotations demonstrate the effectiveness of the pipeline, which can be scaled by adding more 3D assets.
@@ 5.2 Implementation Details
Unbounded Layout Generator. The codebook size $d_{K}$ is set to $512$, with each code having a dimension $d_{C}$ of $512$. Height field and semantic map patches are cropped to 512$\times$512 and compressed by a factor of $16$. The loss weights are $\lambda_{\rm R}=10$, $\lambda_{\rm S}=10$, and $\lambda_{\rm E}=1$.
@@
Stuff and Instance Generators. The City Background Generator, Building Instance Generator, and Vehicle Instance Generator are trained with an Adam optimizer, using $\beta=(0,0.999)$ and a learning rate of $10^{-4}$. The discriminators use the same optimizer settings with a learning rate of $10^{-5}$. Training runs for 298,500 iterations with a batch size of $8$, and images are randomly cropped to 192$\times$192 resolution.
@@ 5.4 Ablation Studies
TABLE III: Quantitative Comparison of Ubounded Layout Generator (UIG). The best values are highlighted in bold. The generated images are centrally cropped to a size of 4096$\times$4096.
@@ G L Hash SinCos FID $\downarrow$ KID $\downarrow$ DE $\downarrow$ CE $\downarrow$
Effectiveness of Building Instance Generator. We highlight the essential role of Building Instance Generator (BIG) in achieving successful unbounded 4D city generation. To validate its effectiveness, we perform an ablation study for BIG. We first compare BIG with two alternative designs: (1) Removing BIG from CityDreamer4D, effectively reverting the model to SceneDreamer, and (2) Generating all buildings simultaneously using BIG without incorporating instance labels. As shown in the first two rows of Table IV and Fig. 8(a)-(b), both alternative designs result in significant degradation in generation quality, underscoring the importance of BIG and instance labels. Scene parameterization directly impacts the quality of 4D city generation. BIG uses vanilla SinCos positional encoding with pixel-wise features from the local encoder. To demonstrate the effectiveness of the scene parameterization in BIG, we compare BIG with other alternative scene parameterization designs.
@@
Scene parameterization plays a critical role in the quality of 4D city generation. BIG leverages vanilla SinCos positional encoding combined with pixel-wise features from the local encoder. To evaluate the effectiveness of BIG’s scene parameterization, we compare it with alternative designs. As shown in the last four rows of Table IV and Fig. 8(c)-(f), using generative hash grid positional encoding results in distorted building façades, while Global Encoders with SinCos encoding introduce repetitive façade patterns. These comparisons emphasize the significance of BIG’s well-designed parameterization in achieving realistic and varied results.
Effectiveness of Vehicle Instance Generator. Vehicle Instance Generator (VIG) plays a critical role in generating vehicles within 4D cities. To validate its effectiveness, we conduct an ablation study on VIG. We compare it with two alternative designs: (1) Removing VIG from CityDreamer4D and treating vehicles as background stuff, allowing City Background Generator to handle their generation, and (2) Generating vehicles without canonicalization, meaning they are not produced in a canonical feature space. As shown in the first two rows of Table V and Fig. 9(a)-(b), both alternative designs lead to severe distortions in the generated results, highlighting the importance of VIG and canonicalization. Scene parameterization is equally critical in VIG. To validate this, we compare different scene parameterization designs within VIG. Currently, VIG uses vanilla SinCos positional encoding combined with global-level features from the global encoder. In the canonical feature space, combining global-level features with 3D coordinates allows the network to better share features across different vehicles, facilitating better convergence. As shown in the last row of Table V and Fig. 9(f), using a local encoder with SinCos positional encoding, as in BIG, makes learning more challenging, resulting in incomplete vehicle shapes. Similarly, using generative hash grid in VIG leads to structural distortions by complicating the network’s ability to associate texture features with 3D coordinates, as illustrated in Fig. 9(c) and 9(e) as well as the 3rd and 5th rows of Table V.

### v4

@@ 1 Introduction
To address these problems, we propose CityDreamer4D, a compositional generative model designed for unbounded 4D cities. As shown in Fig. 1, the unbounded 4D city generation framework separates dynamic objects from static scenes. Static scenes are defined by the city layout from Unbounded Layout Generator, arranging elements like roads, highways, vegetation, and buildings, with the capability to extrapolate to unbounded areas. Dynamic objects, such as vehicles, are defined by traffic scenarios generated by Traffic Scenario Generator, which determines their spatial positioning on high-definition (HD) maps derived from city layouts. Unlike existing methods that use a single module for all objects, CityDreamer4D divides the generation process into three distinct modules: Building Instance Generator for buildings, Vehicle Instance Generator for vehicles, and City Background Generator for background stuff. These generators leverage a highly compact bird’s-eye-view (BEV) scene representation to ensure efficiency and scalability. The scene parameterization is designed to address the unique characteristics of background stuff and instances: background stuff often features similar appearances with irregular textures, while buildings and vehicles display diverse appearances with regular periodic patterns. To handle these variations, we use generative hash grids for the background and apply periodic positional encodings to each instance. We also place buildings in an object-centric coordinate space and vehicles in an object-canonical coordinate space, using specialized methods designed to capture their compact shapes. Compositor combines the rendered background stuff with the building and vehicle instances to create a unified image.
@@
A preliminary version of this work, named CityDreamer, has been published in CVPR 2024 [34]. We make several extensions in this work compared to the preliminary version. 1) We evolve CityDreamer into CityDreamer4D, enabling 4D city generation through Traffic Scenario Generator and Vehicle Instance Generator, effectively separating dynamic objects from static scenes. 2) We enhance the highly compact BEV representation by incorporating an additional bottom-up height map, enabling the representation of hollow structures in cities, such as highways. 3) We propose Traffic Scenario Generator, which creates HD maps from city layouts to produce realistic traffic scenarios with vehicles in unbounded cities. 4) We introduce Vehicle Instance Generator, designed to generate vehicle instances within cities using a novel scene parameterization method grounded on the canonical feature space. 5) We build the CityTopia dataset, offering nearly 40k high-quality street-view and drone-view images with both 2D and 3D semantic and instance annotations.
@@ 2 Related Works
Unlike advanced 2D generative models that mainly focus on individual categories or familiar objects, generating scene-level content is more challenging due to the vast diversity and complexity of scenes [62]. Earlier methods [15, 16] generate scenes by synthesizing videos, but they lack 3D awareness and fail to ensure 3D consistency. Semantic image synthesis approaches [24, 63] have shown promising results in generating scene-level content by conditioning on pixel-wise dense correspondences, like semantic segmentation or depth maps. Several techniques [18, 19, 20] generate 3D scenes by performing inpainting and outpainting on RGB images or feature maps, though most can only interpolate or extrapolate a limited distance from the input views and lack true generative capabilities. Significant research has investigated procedural content generation (PCG) for creating natural [64, 21], indoor [65, 66, 67], and urban scenes [68, 69, 23, 70], but the diversity of the generated scenes is limited by the 3D assets used. Recent methods [26, 7, 9] achieve 3D-consistent scenes at an infinite scale through unbounded layout extrapolation. Other approaches [71, 72, 73] focus on indoor scene synthesis, relying on costly 3D datasets [74, 75] or CAD object retrieval [76, 77, 78].
Figure 1: Overview of CityDreamer4D. 4D city generation comprises static and dynamic scenes, conditioned on city layout $\mathbf{L}$ and time-varying traffic scenario $\mathbf{T}_{t}$, generated by the Unbounded Layout and Traffic Scenario Generators, respectively. City Background Generator uses $\mathbf{L}$ to create background images $\mathbf{\hat{I}}_{G}$ for stuff like roads, vegetation, and the sky, while Building Instance Generator renders the buildings $\{\mathbf{\hat{I}}_{B_{i}}\}$ within the city. Using $\mathbf{T}_{t}$, Vehicle Instance Generator generates vehicles $\{\mathbf{\hat{I}}_{V_{i}}^{t}\}$ at time step $t$. Finally, Compositor combines the rendered background, buildings, and vehicles into a unified and coherent image $\mathbf{\hat{I}}_{C}^{t}$. “Gen.”, “Mod.“, “Cond.”, “BG.”, “BLDG.”, and “VEH.” denote “Generation”, “Modulation”, “Condition”, “Background”, “Building”, and “Vehicle”, respectively.
@@ 3.1 Unbounded Layout Generator
where 0 denotes empty spaces in the 3D volumes.
@@ 3.2 Traffic Scenario Generator
where 0 denotes empty spaces in the 3D volumes.
Traffic Scenario Generation. The generation of traffic scenario $\mathcal{T}$ is conceptualized as the frame-by-frame production of semantic maps $\mathbf{S}_{T}=\left\{\mathbf{S}_{T_{t}}\right\}_{t=1}^{n^{T}}$ and height fields $\mathbf{H}_{T}=\left\{\mathbf{H}_{T_{t}}\right\}_{t=1}^{n^{T}}$. To guarantee realistic and continuous placement of dynamic objects, a high-definition (HD) map is derived from the city layout $\mathbf{L}$. Unlike the city layout, which only specifies the positions of roads and highways, the HD map includes details about lanes, intersections, and traffic signals. Using the generated HD map, an off-the-shelf model [87] determines the per-frame bounding boxes of dynamic objects. The corresponding semantic map and height field are generated based on the bounding boxes.
@@ 3.3 City Background Generator
$C({\mathbf{r}})=\int^{\infty}_{0}A(t)\mathbf{c}(\mathbf{f}_{G}^{\mathbf{r}(t)},l(\mathbf{r}(t)))\boldsymbol{\sigma}(\mathbf{f}_{G}^{\mathbf{r}(t)})dt$ (6)
@@ 3.4 Building Instance Generator
Scene Representation. Building Instance Generator also employs the BEV scene representation. It extracts a local window $\mathbf{L}^{B_{i}}$ from the city layout $\mathbf{L}$ with dimensions $N_{B}^{H}\times N_{B}^{W}\times N_{B}^{D}$. This window is centered around the 2D coordinates $(c_{x}^{B_{i}},c_{y}^{B_{i}})$ of the building instance $B_{i}$. The height field and semantic map used to construct $\mathbf{L}^{B_{i}}$ are represented as $\mathbf{H}_{L}^{B_{i}}$ and $\mathbf{S}_{L}^{B_{i}}$, respectively. Since all buildings share the same semantic label in $\mathbf{S}_{L}$, we perform building instantiation by detecting connected components. Notably, real-world building facades and roofs exhibit distinct visual distributions. To capture this, we assign different semantic labels to the facade and roof of each building instance $\rm B_{i}$ in $\mathbf{L}^{B_{i}}$, with the roof assigned to the top-most voxel layer. All other building instances are excluded from $\mathbf{L}^{B_{i}}$ by assigning them a value of 0.
@@
Note that $\mathcal{O}(\cdot)$ is applied separately to each element of the feature $\mathbf{x}$, with the values normalized to the range [1,1][-1,1].
@@
$C({\mathbf{r}})=\int^{\infty}_{0}A(t)\mathbf{c}(\mathbf{f}_{B_{i}}^{\mathbf{r}(t)},\mathbf{z},l(\mathbf{r}(t)))\boldsymbol{\sigma}(\mathbf{f}_{B_{i}}^{\mathbf{r}(t)})dt$ (11)
@@
($\times 10^{3}$) (km2) Day Night Street Aerial 2D Sem. 2D Inst. 3D Sem. 3D Inst.
@@
CityTopia 37.5 11 36 Synthetic Voxel
@@ 3.5 Vehical Instance Generator
Scene Representation. Vehicle Instance Generator, like Building Instance Generator, leverages the BEV scene representation. It extracts a local window $\mathbf{T}_{t}^{V_{i}}$ from the traffic scenario $\mathbf{T}_{t}$, with dimensions $N_{V}^{H}\times N_{V}^{W}\times N_{V}^{D}$, to generate the vehicle instances within the scene. This window is centered around the 2D coordinates $(c_{x}^{V_{i}},c_{y}^{V_{i}})$ of the vehicle instance $V_{i}$. The height field and semantic map used to construct $\mathbf{T}_{t}^{V_{i}}$ are represented as $\mathbf{H}_{T_{t}}^{V_{i}}$ and $\mathbf{S}_{T_{t}}^{V_{i}}$, respectively. Unlike buildings, vehicle instances are instantiated during the generation of the traffic scenario. Instances other than $V_{i}$ are removed from $\mathbf{T}_{t}^{V_{i}}$ by assigning them a value of 0.
@@
where $\theta\in(-180\degree,180\degree]$ denotes the yaw angle, indicating the vehicle’s heading in the XY-plane relative to the y-y-axis, while $\gamma\in(-90\degree,90\degree)$ represents the pitch angle, with positive or negative values indicating upward or downward tilt relative to the XY-plane. The feature $\mathbf{f}_{\rm V_{i}}^{(\mathbf{p}^{C},t)}$ corresponding to the vehicle $V_{i}$ at time step $t$ for $\mathbf{p}^{C}$ is derived as
@@ 3.6 Compositor
$\mathbf{I}_{C}^{t}=\mathbf{\hat{I}}_{G}\mathbf{M}_{G}+\sum_{i=1}^{n_{B}}{\mathbf{\hat{I}}_{B_{i}}\mathbf{M}_{B_{i}}}+\sum_{i=1}^{n_{V}}{\mathbf{\hat{I}}_{V_{i}}^{t}\mathbf{M}_{V_{i}}^{t}}$ (18)
@@ 4.3 CityTopia Dataset
Virtual City Generation. To build the CityTopia dataset, we design 11 virtual cities in Houdini and Unreal Engine, generating 3D annotations and realistic images with controlled lighting conditions. As illustrated in Fig. 3a, we use a diverse, high-quality set of approximately 5,000 3D assets from the CitySample project [33] to procedurally generate a city prototype in Houdini [^1] https://www.sidefx.com. This city prototype stores the 6D poses of all 3D assets within the city. Through surface sampling, we can assign each 3D point a semantic and instance label, and by instantiating the city prototype in Unreal Engine [^2] https://www.unrealengine.com, we produce a fully generated virtual city.
Image Collection. Once the virtual city is instantiated in Unreal Engine, camera trajectories are set to generate 3,000 images for cities with buildings and 7,500 for a vehicle-only city. Daytime and nighttime scenes are rendered for each trajectory, with sunlight removed to help the network more easily learn lighting consistency during the generation process. To avoid Moiré effects, each image is sampled 8x spatially and 32x temporally during rendering. As shown in Fig. 3c, the CityTopia dataset provides a wider range of viewpoints, shown by its broader elevation angles compared to the GoogleEarth dataset, as well as more street-level perspectives, evidenced by the large number of images taken at near-zero altitude.
2D and 3D Annotation. Since the precise 3D annotations are natively generated from the virtual city pipeline, once the camera poses are set in Unreal Engine, 2D annotations are produced by projecting the 3D annotations using the given camera poses. Fig. 3b highlights the perfect alignment of 2D and 3D instance annotations with both street-view and aerial-view images in the CityTopia dataset. The last row features a vehicle-only scene, enhancing vehicle generation learning. The accurate vehicle annotations demonstrate the effectiveness of the pipeline, which can be scaled by adding more 3D assets.
@@ 5.2 Implementation Details
Unbounded Layout Generator. The codebook size $d_{K}$ is set to $512$, with each code having a dimension $d_{C}$ of $512$. Height field and semantic map patches are cropped to $\times512$ and compressed by a factor of $16$. The loss weights are $\lambda_{\rm R}=10$, $\lambda_{\rm S}=10$, and $\lambda_{\rm E}=1$.
@@
Stuff and Instance Generators. The City Background Generator, Building Instance Generator, and Vehicle Instance Generator are trained with an Adam optimizer, using $\beta=(0,0.999)$ and a learning rate of $10^{-4}$. The discriminators use the same optimizer settings with a learning rate of $10^{-5}$. Training runs for 298,500 iterations with a batch size of $8$, and images are randomly cropped to $\times192$ resolution.
@@ 5.4 Ablation Studies
TABLE III: Quantitative Comparison of Unbounded Layout Generator (ULG). The best values are highlighted in bold. The generated images are centrally cropped to a size of $\times4096$.
@@ G L Hash SinCos FID $\downarrow$ KID $\downarrow$ DE $\downarrow$ CE $\downarrow$
Effectiveness of Building Instance Generator. We highlight the essential role of Building Instance Generator (BIG) in achieving successful unbounded 4D city generation. To validate its effectiveness, we perform an ablation study for BIG. We first compare BIG with two alternative designs: (1) Removing BIG from CityDreamer4D, effectively reverting the model to SceneDreamer, and (2) Generating all buildings simultaneously using BIG without incorporating instance labels. As shown in the first two rows of Table IV and Fig. 8a-b, both alternative designs result in significant degradation in generation quality, underscoring the importance of BIG and instance labels. Scene parameterization directly impacts the quality of 4D city generation. BIG uses vanilla SinCos positional encoding with pixel-wise features from the local encoder. To demonstrate the effectiveness of the scene parameterization in BIG, we compare BIG with other alternative scene parameterization designs.
@@
Scene parameterization plays a critical role in the quality of 4D city generation. BIG leverages vanilla SinCos positional encoding combined with pixel-wise features from the local encoder. To evaluate the effectiveness of BIG’s scene parameterization, we compare it with alternative designs. As shown in the last four rows of Table IV and Fig. 8c-f, using generative hash grid positional encoding results in distorted building façades, while Global Encoders with SinCos encoding introduce repetitive façade patterns. These comparisons emphasize the significance of BIG’s well-designed parameterization in achieving realistic and varied results.
Effectiveness of Vehicle Instance Generator. Vehicle Instance Generator (VIG) plays a critical role in generating vehicles within 4D cities. To validate its effectiveness, we conduct an ablation study on VIG. We compare it with two alternative designs: (1) Removing VIG from CityDreamer4D and treating vehicles as background stuff, allowing City Background Generator to handle their generation, and (2) Generating vehicles without canonicalization, meaning they are not produced in a canonical feature space. As shown in the first two rows of Table V and Fig. 9a-b, both alternative designs lead to severe distortions in the generated results, highlighting the importance of VIG and canonicalization. Scene parameterization is equally critical in VIG. To validate this, we compare different scene parameterization designs within VIG. Currently, VIG uses vanilla SinCos positional encoding combined with global-level features from the global encoder. In the canonical feature space, combining global-level features with 3D coordinates allows the network to better share features across different vehicles, facilitating better convergence. As shown in the last row of Table V and Fig. 9f, using a local encoder with SinCos positional encoding, as in BIG, makes learning more challenging, resulting in incomplete vehicle shapes. Similarly, using generative hash grid in VIG leads to structural distortions by complicating the network’s ability to associate texture features with 3D coordinates, as illustrated in Fig. 9c and 9e as well as the 3rd and 5th rows of Table V.